OpenStudy (anonymous):

Suppose that S varies directly as the 4/5 power of T, and that S=64 when T=32. Find S when T=243

OpenStudy (kropot72):

\[32=2^{5}\] So substituting the above value for 32 we get \[64=k \times (2^{5})^{\frac{4}{5}}=k \times 2^{4}.......(1)\] Now you can solve equation (1) to find the value of k, the constant of proportionality. Also \[243=3^{5}\] Therefore \[S=k \times (3^{5})^{\frac{4}{5}}=k \times 3^{4}..........(2)\] The solution can be found by plugging the value for k into equation (2)